OpenStudy (anonymous):

m<=60-2x and m<5=70-4x a.50 b.10 c.5 d.4 Solve for x

OpenStudy (anonymous):

@pgpilot326 this what Im doing ..

OpenStudy (anonymous):

what is the 2nd part? m < 5 = 70-4x? if so, then 5 = 70-4x. use this to solve for x

OpenStudy (anonymous):

ain't no 2nd part ?

OpenStudy (anonymous):

after the "and"

OpenStudy (anonymous):

m<1=60-2x and m<5 =70-4x ..

OpenStudy (anonymous):

@pgpilot326 even if we solve 5 = 70 - 4x, we will get 65/4 = x which is not available in options

OpenStudy (anonymous):

the problem statement is confusing... what are you asking?

OpenStudy (anonymous):

can you send a pic or screen shot of the actual problem?

OpenStudy (anonymous):

m<1=60-2x and m<5 =70-4x a.50 b.10 c.5 d.4 i can try to send you one